if the perimeter of a rectangle is 460 ft the length of the rectangle is 30 more than the wight. how to you find the wight?
Answer by PRMath(133)

The perimeter is the measure all around a shape. In other words, for a rectangle, you would add the following:

Length + Length + Width + Width = Perimeter. You can rewrite that with "L" for length and "W" for width, which will be:

L + L + W + W = P

You are told that the length ("L") is 30 more than the width ("W").

So, the Length is 30 more than the width, which translated says:
The Length: L
is: =
30 more than the width: 30 + W

This is: L = 30 + W

Let's use L + L + W + W = P and everywhere you have an "L", let's replace it with what we know about "L" which is L = 30 + W.

Also, we will plug in what we know about "P" (perimeter). We know the perimeter is 460.

L + L + W + W = P (perimeter equation)
30 + W + 30 + W + W + W = 460 (see where we put in 30 + W everywhere there was an "L" variable?)
60 + 4W = 460 (see where we combined like terms?)
4W = 400 (subtract 60 from both sides to start to isolate the W)
W = 100 (divide both sides by 4 to isolate the W)

Now we know that the width is 100.
If the width is 100, then the length is 30 more than the width or: 100 + 30 which is 130.

Does this all work out? Let's go back to our original equation and plug in what we know:


L + L + W + W = P
130 + 130 + 100 + 100 = 460

Our info checks out. :-)

I hope this helps you.